OpenStudy (anonymous):

(1, 0) is an ordered pair of the function f(x) = 1 + x. True False

OpenStudy (anonymous):

false. If you put 1 into the equation, because it is the x value of the pair, then 1+x=2.

geerky42 (geerky42):

input is 1, output is 0. let start with input: \(f(1)=1+1=2\) 2 is output, so false.

OpenStudy (anonymous):

thank u sooo much!!

OpenStudy (anonymous):

both of u actually